OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -3, 5 of multiplicity 2. Explain how you would verify the zeros of f(x)

OpenStudy (whpalmer4):

A polynomial with roots a, b, c, d can be written as \[f(x) = (x-a)(x-b)(x-c)(x-d) = 0\]We know the roots are -3 and 5, two of each, so that means we have a = b = -3 and c = d = 5, or \[f(x) = (x+3)(x+3)(x-5)(x-5) = 0\] Multiply that sucker out! To verify the zeros, you just plug those values of x in and make sure that the result is 0.

OpenStudy (anonymous):

do i have to multiply to find the answer? @whpalmer4

OpenStudy (whpalmer4):

Hey, why did you re-ask this question after I answered it?